OpenStudy (anonymous):

7. Determine three consecutive odd integers if the square of the largest integer is 33 less than the sum of the squares of the two smaller integers.

OpenStudy (mertsj):

x=first odd integer x+2 = next odd integer x+4 = third odd integer (x+4)^2=x^2+(x+2)^2-33

OpenStudy (anonymous):

thanks! thats exactly what i needed to get me started

OpenStudy (anonymous):

after factoring i've got x=9 and x=-5, what am i supposed to do next?

OpenStudy (anonymous):

wich x do i use for (x+4), x and (x+2)

OpenStudy (anonymous):

*which

OpenStudy (mertsj):

You have two sets of answers: 9, 11, 13 and -5, -3, -1

OpenStudy (anonymous):

yes i got that answer now